# Causal Inference with Missing Exposures, Missing Outcomes, and Dependence

### v1

@@ 3.4 Missing Exposures and Outcomes at Baseline and Follow-up
To define the causal target parameter in this setting, we first consider the counterfactual outcome at baseline under hypothetical interventions to ensure exposure measurement, “set” the exposure level to $A=a$, and ensure outcome measurement at baseline: $Y_{0}^{*}=Y_{0}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1}$. Additionally, we consider the counterfactual outcome at follow-up under the prior interventions as well as a hypothetical intervention to ensure outcome measurement at follow-up among those at risk at baseline: if $Y_{0}=0$, set $\Delta_{Y_{1}}=1$; else set $\Delta_{Y_{1}}=0$. Thereby, this final intervention is a dynamic or personalized one (e.g., [21, 34, 35, 36]). Denote the resulting counterfactual outcome as $Y_{1}^{*}=Y_{1}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1,\Delta_{Y_{1}}=1}$.
@@
Now we can precisely define the causal parameter in terms of the following conditional probability, which captures the counterfactual incidence of the outcome among those at risk at baseline: $\mathbb{P}(Y_{1}^{*}=1\mid Y_{0}^{*}=0)$. Due to conditioning on a counterfactual variable, such parameters are sometimes called “Counterfactual Strata Effects” and have been used in several real-data analyses [21, 5, 22, 23, 24, 25, 26]. These effects are different from Principal Strata Effects, which are defined within subgroups of latent classes that are fundamentally not observable [37, 38, 39, 40]. For example, in SEARCH-TB, principal stratification could be applied to define the effect of alcohol use on incident TB infection among the subset of participants who would have always tested regardless of their alcohol use. Instead, our interest is the effect of alcohol use on incident TB among the entire population of persons at risk.
@@
Then given the observed data $O=(L_{0},\Delta_{A},A,\Delta_{Y_{0}},Y_{0},L_{1},\Delta_{Y_{1}},Y_{1})$, we can identify denominator and numerator, in turn. The denominator $\mathbb{P}(Y_{0}^{*}=0)$ represents the counterfactual prevalence of not having the outcome at baseline and, thus, being at risk. The causal structure for this parameter is analogous to that of Section 3.2, but with an additional measurement indicator for the outcome. Therefore, under analogous assumptions, we can identify $\mathbb{E}(Y_{0}^{*})=\mathbb{E}\big{[}\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big{]}$ with proof in eAppendix C.1. Since we are interested the counterfactual probability of being at risk at baseline $\mathbb{P}(Y_{0}^{*}=0)$, our statistical estimand for the denominator becomes $1-\mathbb{E}\big{[}\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big{]}$.
In our final causal parameter (Eq. 1), the numerator $\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)$ represents the counterfactual probability of having the outcome at follow-up but not at baseline. For ease of notation, let $Z^{*}=\mathbb{I}(Y_{1}^{*}=1,Y_{0}^{*}=0)$ represent the joint indicator of these two counterfactual values. To identify $\mathbb{E}(Z^{*})=\mathbb{P}(Z^{*}=1)$, we need analogous assumptions as for the denominator together with the following. Among those known to be at risk at baseline ($\Delta_{Y_{0}}=1,Y_{0}=0$) and with measured exposure of interest ($\Delta_{A}=1,A=a$):
  • the baseline and time-varying covariates capture the common causes of the joint outcome and follow-up measurement: $Z^{*}\perp\Delta_{Y_{1}}\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}$.
  • there is a positive probability of follow-up measurement within all possible values of $L_{0}$ and $L_{1}$: $P(\Delta_{Y_{1}}=1\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})>0$ a.e..
Under these assumptions and with proof given in eAppendix C.2, the numerator is identified as
@@
Putting it all together, the statistical estimand with missing exposures, missing outcomes at baseline, and missing outcomes at follow-up is given
@@
3.5 Accounting for Participant Dependence
Finally, we outline an approach to account for dependence between participants within groups or clusters, such as households, schools, hospitals, or communities. Such dependence could arise due to shared exposures and/or the spread of social behaviors or infectious diseases. In our running example, alcohol use may be influenced by members of a participant’s social circle, and TB is transmitted from person to person. This dependence should be reflected in the corresponding causal model (e.g., [41, 42, 43, 44]). Importantly, by following the Causal Roadmap or a similar framework for causal inference [45, 46], we specify causal models encoding our knowledge about the hierarchical data generating process without imposing parametric modeling assumptions — in contrast to more traditional approaches, such as generalizing estimating equations or mixed effects models (e.g., [47, 44, 22, 23, 48]).
Suppose it is reasonable to assume that participants are dependent within households, but households are effectively independent. (We relax this assumption below.) Then our causal model would be specified at the household-level, and identification would consider the influence of other household members as well as community-level factors. Concretely, this may involve including community indicators in $L_{0}$ and summary measures of household-level covariates in $L_{0}$ and $L_{1}$. In other settings, we may need to expand out the causal model and identification results to accommodate arbitrary dependence across a participant’s social network or across participants within a community. The exact form of the causal model and identification result will depend on the application. Going forward, we use “cluster” to refer to any group considered to be the (conditionally) independent unit [47, 23, 48].
4 Statistical Estimation and Inference
@@
In the previous section, we introduced a series of causal models and identification results of increasing complexity. For the resulting statistical estimands, we could use a singly robust estimation approach, such as G-computation or inverse probability weighting (IPW) [12, 11]. Here, we highlight the use of targeted minimum loss-based estimation (TMLE), which is a doubly robust estimation procedure and asymptotically efficient under certain conditions [14]. In TMLE, initial estimates of the relevant pieces of the observed data distribution are updated to achieve the optimal bias-variance trade-off for the estimand and to solve the efficient influence curve equation. Initial estimates are often computed via Super Learner, an ensemble machine learning algorithm using V-fold cross-validation to select an optimal weighted linear combination of predictions from a library of candidate learners [49]. Thereby, TMLE harnesses machine learning to avoid introducing new modeling assumptions during estimation, while supporting valid statistical inference under reasonable conditions. Notably, for ratio-type estimands corresponding to Counterfactual Strata Effects (Eq. 2), we would implement a separate TMLE for the estimand in the numerator (the joint probability) and the estimand in the denominator (the marginal baseline probability) before combining the results.
@@
TMLE is an asymptotically linear estimator and is normally distributed in the large data limit [14]. When the $N$ participants are i.i.d., the estimator minus the estimand behaves like a sample mean in the first order: $\hat{\Psi}-\Psi=\frac{1}{N}\sum_{i=1}^{N}D_{i}+o_{P}(N^{-1/2})$ where $D_{i}$ is the influence curve for participant $i=\{1,\ldots,N\}$ and $o_{P}(N^{-1/2})$ is a second-order remainder term going to zero in probability [50]. The estimated influence curve is used to calculate standard errors, Wald-type confidence intervals, and p-values. Concretely, a 95% confidence interval is constructed using $\hat{\Psi}\pm z_{0.0975}\frac{\hat{\sigma}}{\sqrt{n}}$ where $z_{0.975}$ is the critical value at the 97.5th-percentile of the standard normal and $\hat{\sigma}$ is the standard deviation of the estimated influence curve. For ratio-type estimands (Eq. 2), once the influence curves for the numerator and denominator have been estimated, the Delta method provides an estimate of the influence curve for our overall estimand (i.e., numerator/denominator). Then to calculate measures of association on the difference, ratio, or odds ratio scale, we apply the Delta method a second time to get inference for these types of functionals.
5 Application to SEARCH-TB
@@
We now return to our motivating question: what is the effect of alcohol use on incident TB infection among adults in rural Eastern Uganda? With our multinational and interdisciplinary team, we worked through the Causal Roadmap to specify the Statistical Analysis Plan, including the causal model and the adjustment sets needed to account for confounding, missingness, and participant dependence [45, 52, 53, 31]. Our adjustment set included the SEARCH trial arm, community indicators, household HIV status, as well as individual-level age, sex, and mobility measures. For the primary analysis, we used TMLE with Super Learner to combine estimates from generalized linear models, multivariate adaptive regression splines, and the mean. We conducted influence curve-based inference, accounting for clustering by houshold [47, 23]. In secondary analyses, we considered communities, instead of households, to be the independent unit. To examine the impact of modeling assumptions, we also implemented the inverse probability weighting estimator (IPW) with the same adjustment set, but using parametric regressions to estimate the weights. Unlike TMLE, IPW is singly robust — relying on consistent estimation of the propensity score regressions — and tends to be inefficient [14]. Finally, to examine the impact of our missing data assumptions, we implemented the complete case analysis: the simple ratio of mean outcomes after restricting to participants whose exposure and outcomes were measured. For statistical inference, both IPW and the complete case approach accounted for clustering by household.
@@
In the primary analysis (using TMLE with clustering by household), we found that alcohol use was associated with a 49% increase in the risk of incident TB, after flexibly accounting for confounding and missing data: risk ratio (RR)=1.49 (95%CI: 1.39-1.59) [31]. As shown in Figure 5, secondary analyses with the community as the independent unit yielded very similar results, despite meaningfully reducing the effective sample size from 1,380 households to 9 communities: RR=1.49 (95%CI: 1.37-1.62). In contrast, IPW relying on parametric assumptions resulted in a smaller association and very wide confidence intervals: RR=1.13 (95%CI: 1.00-1.27). Finally, after restricting to participants who responded to questions about alcohol use, tested negative at baseline, and tested again at follow-up, the complete case analysis was the least precise and resulted in very wide confidence intervals: RR=1.18 (95%CI: 0.89-1.57).
@@
Figure 5: Results from SEARCH-TB for the association of alcohol use on incident tuberculosis (TB) infection: “Primary” with TMLE and clustering by household, “Secondary” with TMLE and clustering by community, “IPW” with inverse probability weighting, and “Complete Case” with an unadjusted estimator after excluding participants with missing exposures and outcomes.
6 Discussion
@@
We presented causal models, causal parameters, and identification results for a series of observational studies with increasing levels of missingness. For estimation and inference, we highlighted the use of TMLE with Super Learner to robustly and efficiently estimate the corresponding statistical estimands. Application to real-data from SEARCH-TB demonstrated the real-world consequences of our work. Using TMLE to flexibly account for confounding, missingness, and dependence, we found a 49% relative increase in the risk of incident TB associated with drinking alcohol: RR=1.49 (95%CI: 1.39-1.59). With the same adjustment approach but using parametric regressions, IPW resulted in a smaller association and meaningfully wider confidence intervals: RR=1.13 (95%CI: 1.00-1.27). Finally, without adjustment, the complete case analysis yielded a null association (RR=1.18, 95%CI: 0.89-1.57).
@@
There are limitations to our work. First, we did not provide an exhaustive set of causal models and identification results for all possible studies; however, our approach is generalizable and covers many scenarios arising in public health. Second, in our real-data application, we did not include multiple imputation, which is a common approach for missing data and can also leverage machine learning [54, 55, 56]. Future work is needed to investigate the assumptions, implementation, and performance of MI in settings mirroring our motivating example: (1) missingness on the exposure of interest, (2) missingness on the baseline outcome, crucial to defining the target population, (3) missingness in the final outcome, (4) confounding, and (5) dependence among study participants. Third, we relied on various versions of the missing-at-random assumption throughout. In practice, data may be missing as a result of unobserved variables. When data are missing-not-at-random, we may need to collect additional data and conduct sensitivity/bias analyses [57, 10]. Nonetheless, even when a “causal gap” remains, we have a framework to define a statistical estimand, which is aligned with our research question [45, 58, 52, 53]. Finally, we note that interpretation of causal estimands is nuanced when the exposure influences measurement and outcomes (Figures 3 and 4). By considering hypothetical interventions to ensure outcome measurement, we are blocking part of the exposure’s effect. In other words, our causal estimands correspond to a controlled direct effect — not the total effect. In a competing events setting, Young and colleagues have made similar points and defined causal estimands in terms of “separable” direct and indirect effects [15, 59, 60].
Overall, our goal was to contribute to the missing data literature by providing a framework to avoid complete case analyses and, instead, transparently state the missing data assumptions and robustly estimate the corresponding statistical estimands. To do so, we offered guidance for defining and identifying causal effects in real-world studies with missing exposures, missing outcomes, and dependence. We also demonstrated the real-world consequences of our causal and statistical assumptions in SEARCH-TB.
@@
7 Supplementary Digital Content for “Causal Inference with Missing Exposures, Missing Outcomes, and Dependence”
@@
eAppendix A: Missing exposures (Figure 2 in the main text)
@@
For the corresponding statistical estimand to be well-defined, we also need the following positivity assumptions: $\mathbb{P}(\Delta_{A}=1|L)>0$ a.e. and $\mathbb{P}(A\mid\Delta_{A}=1,L)>0$ a.e..
eAppendix B: Missing Exposures and Outcomes (Figure 3 in the main text)
@@
where the inner expectation averages out the outcome $Y$ given the conditioning set, the middle expectation average out the time-varying covariates $L_{1}$ given the conditioning set, and the outer expectation averages out the baseline covariates $L_{0}$.
For the corresponding statistical estimand to be well-defined, we also need the following positivity assumptions: $\mathbb{P}(\Delta_{Y}=1\mid L_{1},A=a,\Delta_{A}=1,L_{0})>0$ a.e. in addition to the positivity assumption from the previous section.
eAppendix C: Missing Exposures and Outcomes at Baseline and Follow-up (Figure 4 in the main text)
@@
C.1 Identification proof for the denominator
Under the following assumptions, which are analogous to eAppendix A, we can identify 1 minus the denominator:
@@ C.1 Identification proof for the denominator
C.2 Identification proof for the numerator
@@
by $\text{by }Z^{*}=0\text{ when }Y_{0}=1$
@@
For the corresponding statistical estimand to be well-defined, we also need the following positivity assumptions: $P(\Delta_{Y_{1}}=1\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})>0$ a.e. in addition to the positivity assumption for the denominator.

### v2

@@ Introduction
To define the causal target parameter in this setting, we first consider the counterfactual outcome at baseline under hypothetical interventions to ensure exposure measurement, “set” the exposure level to $A=a$, and ensure outcome measurement at baseline: $Y_{0}^{*}=Y_{0}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1}$. Additionally, we consider the counterfactual outcome at endline under the prior interventions as well as a hypothetical intervention to ensure outcome measurement at endline among those known to be at risk at baseline. Concretely, only participants who tested TB-negative at baseline ($\Delta_{Y_{0}}=1$ and $Y_{0}=0$) were approached for re-testing at endline. Thereby, this final intervention is a dynamic or personalized one to set $\Delta_{Y_{1}}$ equal to one if $Y_{0}=0$ and to zero otherwise (e.g., [24, 41, 42, 43]). For simplicity, denote the resulting counterfactual outcome as $Y_{1}^{*}=Y_{1}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1,\Delta_{Y_{1}}=1}$.
@@
Now we can precisely define the causal parameter in terms of the following conditional probability, which captures the counterfactual incidence of the outcome among those at risk at baseline: $\mathbb{P}(Y_{1}^{*}=1\mid Y_{0}^{*}=0)$. Due to conditioning on a counterfactual variable, such parameters are sometimes called “Counterfactual Strata Effects”[24, 5, 25, 26, 27, 28], which are defined by Nakato et al. as “causal estimands where the outcome is only relevant for a group whose membership is subject to missingness and/or impacted by the exposure” [29]. These effects are different from Principal Strata Effects, which are defined within subgroups of latent classes that are fundamentally not observable [44, 45, 46, 47]. For example, in SEARCH-TB, principal stratification could be applied to define the effect of alcohol use on incident TB infection among the subset of participants who would have always tested regardless of their alcohol use. Instead, our interest is the effect of alcohol use on incident TB among the entire population of persons at risk at baseline.
@@
Then given the observed data $O=(L_{0},\Delta_{A},A,\Delta_{Y_{0}},Y_{0},L_{1},\Delta_{Y_{1}},Y_{1})$, we can identify denominator and numerator, in turn. The denominator $\mathbb{P}(Y_{0}^{*}=0)$ represents the counterfactual prevalence of not having the outcome at baseline and, thus, being at risk. The causal structure for this parameter is analogous to that of Section 3.2, but with an additional measurement indicator for the baseline outcome. Therefore, under analogous assumptions, we can identify $\mathbb{E}(Y_{0}^{*})=\mathbb{E}\big[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big]$ with proof in Appendix S2.3. Since we are interested the counterfactual probability of being at risk at baseline $\mathbb{P}(Y_{0}^{*}=0)$, our statistical estimand for the denominator becomes $1-\mathbb{E}\big[\mathbb{E}(Y_{0}\mid\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})\big]$.
In our final causal parameter (Eq. 1), the numerator $\mathbb{P}(Y_{1}^{*}=1,Y_{0}^{*}=0)$ represents the counterfactual probability of having the outcome at endline but not at baseline. For ease of notation, let $Z^{*}=\mathbb{I}(Y_{1}^{*}=1,Y_{0}^{*}=0)$ represent the joint indicator of these two counterfactual values. To identify $\mathbb{E}(Z^{*})=\mathbb{P}(Z^{*}=1)$, we need analogous assumptions as for the denominator together with the following. Among those known to be at risk at baseline ($\Delta_{Y_{0}}=1,Y_{0}=0$) and with measured exposure of interest ($\Delta_{A}=1,A=a$): the baseline and time-varying covariates capture the common causes of the joint outcome and endline measurement: $Z^{*}\perp\Delta_{Y_{1}}\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0}$. We also need there to be a positive probability of follow-up measurement within all possible values of $L_{0}$ and $L_{1}$: $\mathbb{P}(\Delta_{Y_{1}}=1\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})>0$ a.e. Under these assumptions and with proof given in Appendix S2.3, the numerator is identified as
@@
Putting it all together, the statistical estimand with missing exposures, missing outcomes at baseline, and missing outcomes at endline is given by
@@
Statistical Estimation and Inference
@@
In the previous section, we introduced a series of causal models and identification results of increasing complexity. For the resulting statistical estimands, we could use a singly robust estimation approach, such as standardization (a.k.a., “G-computation”) or inverse probability weighting (IPW) [9, 13]. Here, we highlight the use of TMLE, which is a doubly robust estimation procedure and asymptotically efficient under certain conditions [11]. In TMLE, initial estimates of the relevant pieces of the observed data distribution are updated to achieve the optimal bias-variance trade-off for the estimand and to solve the efficient influence curve equation. Initial estimates are often computed via Super Learner, an ensemble machine learning algorithm using V-fold cross-validation to select an optimal weighted linear combination of predictions from a library of candidate learners [48]. Thereby, TMLE leverages machine learning to avoid introducing new modeling assumptions during estimation, while supporting valid statistical inference under reasonable conditions. Notably, for ratio-type estimands corresponding to Counterfactual Strata Effects (Eq. 2), we would implement a separate TMLE for the estimand in the numerator (the joint probability) and the estimand in the denominator (the marginal baseline probability) before combining the results.
@@
TMLE is an asymptotically linear estimator and is normally distributed in the large data limit [11]. The estimator minus the estimand behaves like a sample mean in the first order: $\hat{\Psi}-\Psi=\frac{1}{N}\sum_{i=1}^{N}D_{i}+o_{P}(N^{-1/2})$ where $D_{i}$ is the influence curve for participant $i=\{1,\ldots,N\}$ and $o_{P}(N^{-1/2})$ is a second-order remainder term going to zero in probability [49]. The estimated influence curve is used to calculate standard errors, Wald-type confidence intervals, and p-values. Concretely, a 95% confidence interval is constructed using $\hat{\Psi}\pm z_{0.0975}\frac{\hat{\sigma}}{\sqrt{n}}$ where $z_{0.975}$ is the critical value at the 97.5th-percentile of the standard normal and $\hat{\sigma}$ is the standard deviation of the estimated influence curve. For ratio-type estimands (Eq. 2), once the influence curves for the numerator and denominator have been estimated, the Delta method provides an estimate of the influence curve for our overall estimand. Then to calculate measures of association on the difference, ratio, or odds ratio scale, we apply the Delta method a second time to get inference for these types of functionals.
Application to SEARCH-TB
@@
We now return to our motivating question: what is the effect of alcohol use on incident TB infection among adults in rural Eastern Uganda? With our multinational and interdisciplinary team, we worked through the Causal Roadmap to specify the Statistical Analysis Plan, including the causal model and the adjustment sets [21, 22, 23, 50, 38]. Concretely, the causal model reflected team’s knowledge of study protocol, the study procedures, and the epidemiology of TB in the region. Our adjustment set included the SEARCH trial arm, community indicators, household HIV status, as well as individual-level age, sex, and mobility measures. For the primary analysis, we used TMLE with Super Learner to combine estimates from generalized linear models, multivariate adaptive regression splines, and the mean. We conducted influence curve-based inference, accounting for clustering by household (Appendix S4) [51, 26]. In secondary analyses, we considered communities, instead of households, to be the independent unit. To examine the sensitivity of our results to alternative estimation approaches, we also implemented IPW for the same statistical estimand, but using parametric regressions to estimate the weights. Finally, to examine the impact of handling of our missing data, we took the following “naïve” approach: (1) subset on participants known to be at risk of TB at baseline ($Y_{0}=0$); (2) further subset on participants with measured exposures ($\Delta_{A}=1$) and measured outcomes at endline ($\Delta_{Y_{1}}=1$), and (3) implement TMLE to adjust for confounding and to estimate $\mathbb{E}[\mathbb{E}(Y_{1}\mid A=a,L)\mid Y_{0}=0]$. This approach is inherently flawed, because of the bias induced by inappropriately conditioning on various colliders and mediators (Figure 3). Nonetheless, the approach could be one taken by an analyst aiming to implement a “complete-case” analysis while adjusting for confounding [52, 53, 54]. For statistical inference, both IPW and the naïve approach accounted for clustering by household.
@@
In the primary analysis using TMLE with clustering by household, we found that alcohol use was associated with a 49% increase in the risk of incident TB: risk ratio (RR)=1.49 (95%CI: 1.39-1.59) [38]. As shown in Figure 4, secondary analyses with the community as the independent unit yielded very similar results, despite meaningfully reducing the effective sample size from 1,380 households to 9 communities: RR=1.49 (95%CI: 1.37-1.62); Appendix S4. In contrast, IPW relying on parametric assumptions resulted in a smaller association and confidence intervals overlapping the null: RR=1.13 (95%CI: 1.00-1.27). Finally, after restricting to participants who tested negative at baseline, responded to questions about alcohol use, and tested again at follow-up, the naïve approach was the least precise and resulted in the widest confidence intervals: RR=1.18 (95%CI: 0.89-1.57).
@@
Figure 4: Results from SEARCH-TB for the association of alcohol use on incident tuberculosis (TB) infection: “Primary” with TMLE and clustering by household, “Secondary” with TMLE and clustering by community, “IPW” with inverse probability weighting, and “Naïve” based on subsetting on those at risk at baseline and with measured exposures and outcomes at endline.
Discussion
@@
We presented causal models, causal parameters, and identification results for a series of prospective studies with increasing levels of missingness. For estimation and inference, we highlighted the use of TMLE with Super Learner to robustly and efficiently estimate the corresponding statistical estimands. Application to real-data from SEARCH-TB demonstrated the real-world consequences of our work. Using TMLE to flexibly account for confounding and missingness, we found a 49% relative increase in the risk of incident TB infection associated with drinking alcohol: RR=1.49 (95%CI: 1.39-1.59). For the same statistical estimand but using parametric regressions, IPW resulted in a smaller association and meaningfully wider confidence intervals: RR=1.13 (95%CI: 1.00-1.27). Finally, the naïve approach to condition on the participants with baseline TB risk and measured exposures/outcomes yielded a null association (RR=1.18, 95%CI: 0.89-1.57).
There are several advantages to our framework for handling missing exposures and outcomes when evaluating the causal effects. Unlike the m-DAG approach which aids in identifying the entire joint distribution of the observed data [4], our approach focuses our efforts on the portions of the observed data distribution relevant for the statistical estimand. Furthermore, our approach uses data on all participants, improving efficiency relative to approaches excluding participants with missing data on the relevant variables (sometimes called an “available-case analysis”) [4]. Finally, our approach leads us to statistical estimands that can be robustly and rigorously estimated with modern methods, such as TMLE.
A major limitation to this work is that we did not address missingness on the confounders. Some approaches include mean/median imputation or including missingness indicators in the adjustment set. For further discussion on missing confounders, we refer to the comprehensive work of Williamson et al. [55]. We also focused on cross-sectional or prospective studies. Thus, we did not cover scenarios where the outcome impacts the measurement of other variables. Such scenarios would arise in case-control studies and have been addressed in prior literature (e.g., [56, 57, 58, 59, 54]).
@@
There are other limitations to our work. First, we did not provide an exhaustive set of causal models and identification results for all possible studies; however, our approach is generalizable and covers many scenarios arising in public health. Second, in our real-data demonstration, we did not include multiple imputation, which is a common approach for missing data and can also leverage machine learning [60, 61, 62]. Future work is needed to investigate the assumptions, implementation, and performance of multiple imputation in settings mirroring our motivating example: (1) missingness on the exposure of interest, (2) missingness on the baseline outcome, (3) missingness in the final outcome, (4) confounding, and (5) dependence among study participants. Finally, we relied on various versions of the sequential exchangeability assumption for both confounding and missingness. In practice, data may be missing as a result of unobserved variables, and we may need to collect additional data as well as conduct sensitivity/bias analyses [63, 64]. Nonetheless, even when a “causal gap” remains, we have a framework to define a statistical estimand, which is aligned with our research question [21, 65, 23, 50].
Overall, our goal was to demonstrate how an established framework for estimating causal effects with censored outcomes could be extended to settings with missingness on the exposure, the outcome at baseline, and the outcome at endline. To do so, we provided a series of causal models with increasing complexity and discussed the identification assumptions for each. For estimation and inference, we focused on TMLE, a doubly-robust approach using machine learning and data from all participants (including those with missing exposures or outcomes). Our work was motivated by the investigation of the effect of alcohol use on incident TB infection [38], and our re-analysis demonstrated the real-world consequences of our approach.
@@
Conflicts of Interest: The authors report no conflict of interest in this work.
@@
Supplementary Materials for “Causal Inference with Missing Exposures and Missing Outcomes”
Contents:
  • Appendix S1: More on the classic point-treatment problem
  • Appendix S2: Proofs
  • Appendix S3: Accounting for Outcome Dependence
Appendix S1: More on the classic point-treatment problem
We consider the classic “point-treatment” problem, where we have measured confounding by baseline covariates $L$, a binary exposure $A$ occurring at single time-point, and an outcome $Y$ occurring at the study’s close. This could represent a study of the effect of alcohol use ($A$) on incident TB infection ($Y$) among a representative cohort of persons without TB at baseline. The directed acyclic graph (DAG) and non-parametric structural equation model (NPSEM) for such a study are given in Figure 5A.
Figure 5: Causal models for a classic point-treatment problem with complete measurement of the baseline covariates $L$, the exposure $A$, and the outcome $Y$. On the directed acyclic graph, $U^{*}$ represents unmeasured common causes of at least two variables in $(L,A,Y)$. In panel A provides the causal models under no assumptions about the unmeasured factors. Panels B and C are compatible with the no unmeasured confounders assumption.
To identify our causal target parameter and express it as function of the distribution of the observed data $O=(L,A,Y)$, we would need there to be no unmeasured confounding, which corresponds to the assumption that the baseline covariates $L$ capture all the joint causes of the exposure $A$ and outcome $Y$. This condition is called “the randomization assumption” or “exchangeability”, can be evaluated with the backdoor criterion, and can be represented as $Y^{a}\perp A\mid L$ [9, 10, 11, 12].
Concretely, $\mathbb{E}(Y^{a})$ is not identified in Figure 5A because there is an unmeasured common cause, represented by $U^{*}$, of the exposure $A$ and outcome $Y$. In Figure 5B and C, we show two causal models where this assumption would hold.
Additionally, we need there to be a non-zero probability of having the exposure in all possible values of $L$: $\mathbb{P}(A=a\mid L)>0$ a.e. This is a condition on data support and known as the “positivity assumption”. We note that the consistency assumption, stating that the counterfactual outcome $Y^{a}$ equals the observed outcome $Y$ under exposure $A=a$, holds by design when we define counterfactuals through intervention on the causal model. Under the exchangeability and positivity assumptions, our causal target is equal to the statistical estimand given by the G-computation formula: $\mathbb{E}[\mathbb{E}(Y\mid A=a,L)]$ [9]. Even if these assumptions are not reasonable (e.g., Figure 5A reflects reality), we still have well-defined statistical estimand, on which we can focus our estimation efforts. In other words, we still proceed with estimation and inference $\mathbb{E}[\mathbb{E}(Y\mid A=a,L)]$ , while appropriately accounting for lack of identification in our interpretation [21, 22, 23, 66, 65] For ease of presentation, we provide the causal models without dependence between the unmeasured factors in the main text.
Appendix S2: Proofs
@@
Appendix S2.1: Missing exposures (Figure 1C in the main text)
@@
We again note the consistency assumption holds by our definition of counterfactual outcomes as being derived through interventions on the causal model. For the corresponding statistical estimand to be well-defined, we also need the following positivity assumptions: $\mathbb{P}(\Delta_{A}=1|L)>0$ a.e. and $\mathbb{P}(A\mid\Delta_{A}=1,L)>0$ a.e..
Appendix S2.2: Missing Exposures and Outcomes (Figure 2 in the main text)
@@
where the inner expectation averages out the outcome $Y$ given the conditioning set, the middle expectation average out the time-varying covariates $L_{1}$ given the conditioning set, and the outer expectation averages out the baseline covariates $L_{0}$. For the corresponding statistical estimand to be well-defined, we also need the following positivity assumptions: $\mathbb{P}(\Delta_{Y}=1\mid L_{1},A=a,\Delta_{A}=1,L_{0})>0$ a.e. in addition to the positivity assumptions from the previous section.
Appendix S2.3: Missing Exposures and Outcomes at Baseline and Follow-up (Figure 3 in the main text)
@@
Identification proof for the denominator
Under the following assumptions, which are analogous to Appendix S2.1, we can identify 1 minus the denominator:
@@
Identification proof for the numerator
@@
$\text{by }Z^{*}=0\text{ when }Y_{0}=1$
@@
For the corresponding statistical estimand to be well-defined, we also need the following positivity assumptions: $P(\Delta_{Y_{1}}=1\mid L_{1},Y_{0}=0,\Delta_{Y_{0}}=1,A=a,\Delta_{A}=1,L_{0})>0$ a.e. in addition to the positivity assumptions for the denominator.
Appendix S3: Accounting for Outcome Dependence
Here, we outline an approach to account for the dependence of participant outcomes within groups or clusters, such as households, schools, hospitals, or communities. Such dependence could arise due to shared exposures and/or the spread of social behaviors or infectious diseases. In our running example, TB is transmitted from person to person. This dependence should be reflected in the corresponding causal model (e.g., [67, 68, 69, 70]). By following the Causal Roadmap or a similar framework for causal inference [21, 71], we specify causal models encoding our knowledge about the hierarchical data generating process without imposing parametric modeling assumptions — in contrast to more traditional approaches, such as generalizing estimating equations or mixed effects models (e.g., [51, 70, 25, 26, 72]).
Suppose it is reasonable to assume that participant outcomes are dependent within households, but effectively independent between households. Then our causal model would be specified at the household-level, and identification would consider the influence of other household members as well as community-level factors. (See, for example, [70].) Concretely, this may involve including community indicators in $L_{0}$ and summary measures of household-level covariates in $L_{0}$ and $L_{1}$. The exact form of the causal model and identification result will depend on the application. Going forward, we use “cluster” to refer to any group considered to be the (conditionally) independent unit [51, 26, 72].